Solve the problem by setting up and solving an appropriate algebraic equation.
Suppose that the length of a certain rectangle is 4 meters less than four times its width. The perimeter of the rectangle is 52 meters. Find the length and width of the rectangle.

Answer by richwmiller(17219)

p=2L+2W
L=4W-4
now plug in the numbers
p=2L+2W
52=2(4W-4)+2W
52=8W-8+2W
52=10W-8
60=10W
W=6
L=4W-4
L=4(6)-4
L=20
check
p=2L+2W
52=2*20+2*6
52=40+12
52=52
ok